OpenStudy (anonymous):

what is (2+sqrt4)(2-sqrt4) multiplied?

OpenStudy (anonymous):

0 - because the sqrt4=2 so you got (2+2)(2-2)=0

OpenStudy (apoorvk):

\[(a+b)(a-b) = a^2 - b^2\]

OpenStudy (kropot72):

\[\sqrt{4}==+2 or -2\]

OpenStudy (kropot72):

\[(2+\sqrt{4})(2-\sqrt{4})=4-4=0\]

jhonyy9 (jhonyy9):

4-4=0